What is the LCD of -3/14 and 1/10? Is it 70?
140
3
14
70
Thank you!! Mom

Answer by josgarithmetic(39799)

14=2*7;
10=2*5.

Lowest common denominator needs 2, 5, and 7. Each must appear once in a product.
2*5*7=70